### v4

@@ V-A1 Result Analysis
V-B The effectiveness of Global-Module-Aware Knowledge (RQ2)
In this RQ, our goal is to explore the effectiveness of global-aware information in aiding LLMs to reuse global functions in code generation. We aim to identify the optimal top-k value for retrieving candidate functions to achieve the best global-aware configuration.
Specifically, we employ No Global-Aware as the baseline and this involves providing only Local-Aware Knowledge without Global-Aware Knowledge input. Furthermore, to determine the optimal top-k value, we define four variants, namely, k=1, 5, 10, and 15. We run the total five different configurations on the benchmark to obtain code generation results. Subsequently, we assess the generated code using the same method as in sectionĀ V-A, where each result is evaluated for whether it uses global functions.
@@ V-B The effectiveness of Global-Module-Aware Knowledge (RQ2)
No Global-Aware (k = 0) N/A N/A N/A N/A 0.833
k=1 1.757 0.384 0.516 0.440 0.781
k=5 8.154 0.542 0.703 0.612 0.851
k=10 16.031 0.493 0.562 0.526 0.830
k=15 23.592 0.395 0.531 0.453 0.786
V-B1 Result Analysis
TableĀ II presents the results obtained using ā€œNo Global-Awareā€ and four different top-k values for retrieving global functions. The No Global-Aware has both precision and recall as 0 because of lacking of global-aware information as we expected. Consequently, the generated code only utilizes local functions to accomplish tasks, and it resorts to defining new functions for specific logic if local functions are insufficient. When k=1, the model shows some awareness of global functions to a certain extent, underlined by the precision and recall as 0.384 and 0.516, respectively. At k=5, the model’s performance significantly improves, indicating that the model now has a more comprehensive understanding of global-aware information. However, when k=10, the model’s F1 score drops by 14.05%, indicating a substantial reduction in the model’s ability to reuse global functions. As k rises to 15, the effectiveness of the model decreases further. This is because the overdosed amount (between 10 to 20 functions) of functions is retrieved, introducing excessive interference and hindering accurate reusing.
@@ V-B1 Result Analysis
Finding 2: Global-Aware Information significantly enhances the model’s ability to reuse functions in other code files in the repository. The optimal configuration for retrieving global functions is k=5.
V-C The effectiveness of Third-Party-Library-Aware Knowledge (RQ3)
In sectionĀ III-B3, we delve into Third-Party-Library Aware information, which helps the model recognize pre-installed third-party libraries in the code repository. To validate that point, we compare code generation between with and without Third-Party-Library Aware information. Only Local-Aware and Global-Aware information is fed into the model. Then, we run these two different configurations on the benchmark to obtain code generation results. Notably, we introduce the Library Coverage, the percentage of third-party libraries used in the generated code in the third-part-library base, to further elaborate the extent of reusing.
@@ V-C The effectiveness of Third-Party-Library-Aware Knowledge (RQ3)
No 0.727 0.427 0.538 0.713 0.875
Yes 0.957 0.587 0.727 0.828 0.940
V-C1 Result Analysis
TableĀ III demonstrates that after incorporating Third-Party-Library Aware Knowledge, the model exhibits significant improvements in precision, recall, F1 score, and accuracy, with increases of 31.64%, 37.47%, 35.13%, and 16.13%, respectively. This enhancement can be attributed to the fact that, with Third-Party-Library Aware information, the model is suggested to use a specific subset of libraries, narrowing down its choice range and making it more likely to generate desired code. Our finding is further proven by that methods with Third-Party-Library Aware information outperform those without such knowledge in the Library Coverage metric by 7.43%. However, library coverage does not reach 100% due to occasional instances where the model obstinately believes that there are no viable libraries in the third-party repository to implement the code, which could be attributed to model hallucinationsĀ [57, 58, 59].
@@ V-C1 Result Analysis
V-D The Effectiveness of $A^{3}$odGen Framework (RQ4)
In previous sections, we have individually verified the effectiveness of Local-Aware, Global-Aware, and Third-Party-Library-Aware information. We also explored the optimal configurations. In this section, we aim to integrally validate the effectiveness of our $A^{3}$-CodGen Framework from various perspectives. We compare our $A^{3}$-CodGen with three baselines: vanilla ChatGPT, $A^{1}$ Generation (Local-Aware), and $A^{2}$ Generation (Local-Aware and Global-Aware). We evaluate our framework using four metrics: reuse awareness, reuse correctness, library coverage, and lines of code (LOC). The F1 score serves as a comprehensive measure of reuse awareness. Notably, a lower LOC signifies more concise code.
TABLE IV: Comparison of the overall effect of Framework with Ablation Baselines
@@ V-D The Effectiveness of $A^{3}$odGen Framework (RQ4)
ChatGPT
Local-Aware
Local-Aware
+Global-Aware
Local-Aware
+Global-Aware
+3rd-Party-Lib-Aware
Local Reuse Awareness (F1) N/A 0.683 0.664 0.693
Correctness N/A 0.648 0.646 0.633
Global Reuse Awareness (F1) N/A N/A 0.612 0.632
Correctness N/A N/A 0.604 0.601
3rd-Party Library Reuse Awareness (F1) 0.640 0.659 0.538 0.727
Correctness 0.640 0.742 0.743 0.850
Library Coverage 0.786 0.863 0.875 0.940
Avg Line of Code (LOC) 16.308 16.507 12.598 12.804
V-D1 Result Analysis
Upon evaluating the awareness aspect, as shown in TableĀ IV, $A^{3}$-CodGen demonstrates excellence in local reuse, global reuse, and third-party library reuse. This underscores the effective assistance provided by the three categories of repository-aware knowledge in enabling the model to be cognizant of repository information during code generation.
@@ V-D1 Result Analysis
Finding 4: $A^{3}$-CodeGen framework excels in awaring three types of code repository knowledge, resulting in an excellent performance in terms of local reuse, global reuse, and third-party library reuse. Thanks to its outstanding reuse capabilities, our framework can generate concise code that aligns with the code repository’s working environment.
@@ VI Discussion
New Paradigm for Code Generation Tasks. Code generation tasks have traditionally focused on automating code production for specific requirements, with advancements like GitHub CoPilot and ChatGPT mastering this task. However, real-world code generation extends beyond meeting requirements; it requires generating code aligned with project-specific needs and leveraging existing knowledge for improved readability, maintainability, and practicability. Therefore, we hope this work sheds a light on this emerging but significant code generation topic: generating context-aware code that is compatible with the working environment.
@@
Knowledge Engineering for Code Generation. LLMs are sensitive to prompts and well-designed prompts are decisive in obtaining high-quality resultsĀ [60, 61, 62]. Either unorganized or too lengthy prompts are impractical. Instead, it is advisable to extract semantic-enriched knowledge from the code repository to guide the model to directional generate code. However, how to selectively fuse the different types of knowledge presents a challenge, potentially impacting the model’s awareness of different knowledge types. In SectionĀ V-A1, we find that providing diverse information does not necessarily improve model performance. Meanwhile, in sectionĀ V-D, we find that providing the model with global awareness and third-party library information unexpectedly enhances its ability to reuse native functions. This underscores the interaction of diverse code knowledge types and stresses the significance of efficiently mining and utilizing code knowledge in future studies.
Threats to Validity. The threats to validity of this paper lies in two aspects: 1) Subjective Labeling. Our study involves manual labeling, introducing subjectivity in annotations. We mitigate this by engaging two professional developers for labeling, resolving any instances with ambiguous labels to establish a reliable ground truth. The dataset is public for community validationĀ [19]. 2) Limited Dataset Selection. Experiments use 29 meticulously selected Python code repositories, emphasizing high quality indicated by stars and forks, while avoiding data leaks. These repositories collectively represent the Python programming ecosystem. Rigorous sampling ensures a confidence level of 95% or higher, ensuring the credibility of the ground truth. In the future, we’ll maintain this dataset, providing the community with a valuable resource for evaluating repository-aware code generation.

### v5

@@ V-A3 Result Analysis
V-B RQ2: Is Global Module Aware Knowledge effective in assisting LLMs to utilize global module knowledge during code generation, and what is the optimal ā€œkā€ value for retrieving global functions?
V-B1 Motivation
In this RQ, our goal is to explore the effectiveness of global-aware information in aiding LLMs to reuse global functions in code generation. When using global-aware knowledge, a critical issue arises: if too few candidate functions are retrieved, the model may lack the necessary functions to generate code that correctly reuses global functions. Conversely, retrieving too many candidate functions can lead to the presence of irrelevant functions, making it difficult for the model to identify the required ones. Research shows that irrelevant context can decrease model efficiency, increase costs, and potentially compromise performanceĀ [33]. Therefore, we aim to identify the optimal top-k value for retrieving candidate functions to achieve the best global-aware configuration. Additionally, we conduct an ablation study to evaluate the effectiveness of the What-if Code Generator Unit based on the optimal k value.
V-B2 Experiment Methodology
We employ No Global-Aware as the baseline and this involves providing only Local-Aware Knowledge without Global-Aware Knowledge input. To determine the optimal top-k value, we pre-define four variants: k=1, 5, 10, and 15. These values are chosen because larger projects with more complex call relationships require at least one global function. A single retrieved global function may not be sufficient, and some functions may need to call multiple global functions, so we increase the k value to 5, 10, and 15. This selection ensures we cover a range of potential requirements, balancing between too few and too many candidate functions to maintain model performance and relevance. We run the total five different configurations on the benchmark to obtain code generation results. Subsequently, we assess the generated code using the same method as in sectionĀ V-A, where each result is evaluated for whether it uses global functions.
To further evaluate the effectiveness of our What-if Code Generator Unit, we conduct an experiment using the benchmark, focusing on test cases that call global functions (47 cases in total). We compare two retrieval methods:
    FunDes: Uses only function descriptions for text similarity retrieval.
    FunDes+FunCode: Uses function descriptions for text similarity retrieval and the What-If Code Generator to generate intermediate code snippets for code similarity retrieval.
@@ V-B2 Experiment Methodology
No Global-Aware (k = 0) N/A N/A N/A N/A 0.843
k=1 1.757 0.372 0.533 0.438 0.786
k=5 8.154 0.518 0.717 0.601 0.851
k=10 16.031 0.479 0.583 0.526 0.836
k=15 23.592 0.407 0.583 0.479 0.802
V-B3 Result Analysis
TableĀ II presents the results obtained using ā€œNo Global-Awareā€ and four different top-k values for retrieving global functions. The No Global-Aware has both precision and recall as 0 because of lacking of global-aware information as we expected. Consequently, the generated code only utilizes local functions to accomplish tasks, and it resorts to defining new functions for specific logic if local functions are insufficient. When k=1, the model shows some awareness of global functions to a certain extent, underlined by the precision and recall as 0.372 and 0.533, respectively. At k=5, the model’s performance significantly improves, indicating that the model now has a more comprehensive understanding of global-aware information. However, when k=10, the model’s F1 score drops by 12.48%, indicating a substantial reduction in the model’s ability to reuse global functions. As k rises to 15, the effectiveness of the model decreases further. This is because the overdosed amount (between 10 to 20 functions) of functions is retrieved, introducing excessive interference and hindering accurate reusing.
@@ V-B3 Result Analysis
To further evaluate the effectiveness of our What-if Code Generator Unit, we conduct an experiment using the benchmark. The results showed that the FunDes method retrieved 41.5% of global functions, while the FunDes+FunCode method retrieved 60.6%, representing an almost 20% increase. This demonstrates that our What-if Code Generator Unit enhances the Global Function Retrieval Unit by providing intermediate code for code similarity metrics, thus retrieving more relevant global functions.
Finding 2: Global-Aware Information significantly enhances the model’s ability to reuse functions in other code files in the repository. The optimal configuration for retrieving global functions is k=5. The What-if Code Generator Unit helps retrieve more relevant global functions, further enhancing the Global Function Retrieval Unit.
V-C RQ3: Does Third-Party Library Aware Knowledge significantly impact LLMs’ awareness of available third-party libraries during code generation?
V-C1 Motivation
In sectionĀ III-B3, we delve into Third-Party-Library Aware information, which helps the model recognize pre-installed third-party libraries in the code repository. In this section, we aim to evaluate the effectiveness of third-party-library knowledge.
V-C2 Experiment Methodology
To validate this point, we compare code generation between with and without third-party-library aware information. Only local-aware and global-aware information is fed into the model. Then, we run these two different configurations on the benchmark to obtain code generation results. Notably, we introduce the Library Coverage, the percentage of third-party libraries used in the generated code in the third-part-library base, to further elaborate the extent of reusing.
@@ V-C2 Experiment Methodology
No 0.716 0.429 0.536 0.715 0.875
Yes 0.935 0.585 0.720 0.825 0.940
V-C3 Result Analysis
TableĀ III demonstrates that after incorporating Third-Party-Library Aware Knowledge, the model exhibits significant improvements in precision, recall, F1 score, and accuracy, with increases of 30.59%, 36.36%, 34.33%, and 15.38%, respectively. This enhancement can be attributed to the fact that, with Third-Party-Library Aware information, the model is suggested to use a specific subset of libraries, narrowing down its choice range and making it more likely to generate desired code. Our finding is further proven by that methods with Third-Party-Library Aware information outperform those without such knowledge in the Library Coverage metric by 7.43%. However, library coverage does not reach 100% due to occasional instances where the model obstinately believes that there are no viable libraries in the third-party repository to implement the code, which could be attributed to model hallucinationsĀ [73, 74, 75]. To mitigate these hallucinations, one potential approach is to use a reflective approachĀ [51, 76, 77]. If the model uses a non-pre-installed library, we can inform the model that the library is not expected and prompt it to regenerate the code. This process is repeated until the model uses a pre-installed library. Although we have not systematically validated this approach due to the high labeling effort and API token costs, it represents a feasible method to mitigate model hallucinations.
@@ V-C3 Result Analysis
V-D RQ4: What is the overall performance of our $A^{3}$odGen framework?
V-D1 Motivation
In previous sections, we have individually verified the effectiveness of local-aware, global-aware, and third-party-library-aware information. We also explored the optimal configurations. In this section, we aim to integrally validate the effectiveness of our $A^{3}$-CodGen framework from various perspectives. Therefore, we conduct an ablation study to evaluate $A^{3}$-CodGen in terms of local reuse, global reuse, and third-party reuse. Additionally, as more powerful models are released, we aim to compare the performance of $A^{3}$-CodGen across different foundation models.
V-D2 Experiment Methodology
We first compare our $A^{3}$-CodGen with three baselines using the GPT-3.5-Turbo-16k modelĀ [45]: vanilla ChatGPT, $A^{1}$ Generation (Local-Aware), and $A^{2}$ Generation (Local-Aware and Global-Aware). Then, to investigate whether performance improvements are significant with enhanced model capabilities, we use the latest LLM model, GPT-4oĀ [78], as the foundation model for $A^{3}$-CodGen and compare its performance against GPT-3.5-Turbo-16k. We evaluate our framework using four metrics: reuse awareness, reuse correctness, library coverage, and lines of code (LOC). The F1 score serves as a comprehensive measure of reuse awareness. Notably, a lower LOC signifies more concise code.
TABLE IV: Comparison of the overall effect of the framework with ablation baselines using GPT-3.5-Turbo-16k for $A^{1}$-CodGen, $A^{2}$-CodGen, and $A^{3}$-CodGen configurations, and comparing $A^{3}$-CodGen with GPT-3.5-Turbo-16k and GPT-4o.
@@ V-D2 Experiment Methodology
GPT-3.5-Turbo-16k
$A^{1}$-CodGen (GPT-3.5-Turbo-16k) $A^{2}$-CodGen (GPT-3.5-Turbo-16k) $A^{3}$-CodGen (GPT-3.5-Turbo-16k) $A^{3}$-CodGen (GPT-4o)
Local Reuse Awareness (F1) N/A 0.673 0.673 0.694 0.718
Correctness (F1) N/A 0.675 0.645 0.632 0.879
Global Reuse Awareness (F1) N/A N/A 0.601 0.606 0.701
Correctness (F1) N/A N/A 0.611 0.591 0.676
3rd-Party Library Reuse Awareness (F1) 0.635 0.652 0.536 0.720 0.898
Correctness (F1) 0.644 0.734 0.735 0.852 0.894
Library Coverage 0.786 0.863 0.875 0.940 0.976
Avg Line of Code (LOC) 16.308 16.507 12.598 12.804 13.885
V-D3 Result Analysis
Upon evaluating the awareness aspect, as shown in TableĀ IV, $A^{3}$-CodGen based on GPT-3.5-Turbo-16k demonstrates excellence in local reuse, global reuse, and third-party library reuse. This underscores the effective assistance provided by the three categories of repository-aware knowledge in enabling the model to be cognizant of repository information during code generation.
@@ V-D3 Result Analysis
By examining the last two columns (TableĀ IV) of $A^{3}$-CodGen with GPT-3.5-Turbo-16k and GPT-4o, we observe significant improvements in local reuse, global reuse, and third-party library reuse for GPT-4o compared to GPT-3.5-Turbo-16k. This demonstrates that the same prompts achieve better performance as the foundation model’s capabilities improve. Thus, our $A^{3}$-CodGen’s performance enhances with the increased ability of the foundation model. More powerful foundation models in future will lead to even better results. However, the code generated by $A^{3}$-CodGen based on GPT-4o is slightly more complex, with an average LOC increase of one line, compared to GPT-3.5-Turbo-16k. This is because GPT-4o considers more logical coherence between code segments, resulting in slightly more complex code. Nonetheless, the generated code remains more concise than that produced by Vanilla GPT-based code generation.
@@
Finding 4: $A^{3}$-CodeGen framework excels in awaring three types of code repository knowledge, resulting in an excellent performance in terms of local reuse, global reuse, and third-party library reuse. Thanks to its outstanding reuse capabilities, our framework can generate concise code that aligns with the code repository’s working environment. Additionally, $A^{3}$-CodGen’s performance improves with the increased capabilities of the foundation model.
V-E RQ5: How does $A^{3}$odGen framework compare to GitHub Copilot?
V-E1 Motivation
Given advancements in AI-driven code generation, it is essential to compare our $A^{3}$-CodGen framework with leading tools like GitHub CopilotĀ [9]. GitHub Copilot generates code by leveraging local and global contextual information from neighboring tabs, [^2] https://github.blog/ai-and-ml/github-copilot/how-github-copilot-is-getting-better-at-understanding-your-code/ similar to our context-aware mechanisms. If Copilot can outperform our approach, further development of our method may be unnecessary. However, if it cannot, our goal is to identify the limitations in Copilot’s design for collecting context and explore how to enhance its code reuse capabilities using insights from our framework.
V-E2 Experiment Methodology
For each function in the benchmark dataset, we locate the function within the repository using the repository name and fully qualified name, then remove the function code and the corresponding import statements. We provide the function’s comment and definition to trigger GitHub Copilot to generate the code by pressing the ā€˜tab’ key. To assist Copilot in generating third-party library imports, we add a comment #generate import statements if necessary:. If Copilot makes no suggestion, it indicates no import statements are needed. We retain all necessary local functions that could be reused within the function. Finally, we used the same labeling method mentioned in SectionĀ IV-C to label the generated code. The GitHub Copilot plugin version in this experiment is 1.5.15.6094, and we disable the auto-update in settings. We compare $A^{3}$-CodGen with GitHub CoPilot in terms of local reuse, global reuse, and third-party library reuse. We use the RepoEval dataset mentioned in SectionĀ IV-B2.
Figure 5: Comparison between GitHub CoPilot and $A^{3}$-CodGen.
V-E3 Result Analysis
To evaluate the statistical significance of the differences between $A^{3}$-CodGen and CoPilot, we conduct Welch’s t-testĀ [79] with a significance threshold of $p<0.05$. The p-values for local reuse awareness and correctness are 0.0696 and 0.5590, indicating no significant difference. For global reuse awareness and correctness, the p-values are 0.0144 and 1.46e-06, and for third-party library reuse awareness, correctness, and coverage, the p-values are 1.39e-11, 0.0231, and 0.0176, indicating significant differences.
Fig.Ā 5 compares GitHub Copilot and $A^{3}$-CodGen. For local reuse, GitHub CoPilot performs excellently with an F1 score of 0.790 for local reuse awareness, while $A^{3}$-CodGen achieves an F1 score of 0.718. This is because CoPilot reviews all working context details (including function definitions, function code, and descriptions) when generating the code. In terms of local reuse correctness, $A^{3}$-CodGen reaches an F1 score of 0.879, slightly better than GitHub CoPilot’s 0.857, due to our structured approach to local knowledge.
However, for global reuse awareness and correctness, CoPilot’s performance is suboptimal, with F1 scores of 0.447 and 0.136, respectively. CoPilot’s workflow, although capable of utilizing neighboring tabs’ context, struggles with global function calls as necessary global functions might exist in unopened tabs. Opening all repository tabs is impractical, and CoPilot does not structure information for completion, leading to chaotic context. In contrast, $A^{3}$-CodGen retrieves relevant global functions and structures them, facilitating easier and more accurate reuse. For third-party library reuse, $A^{3}$-CodGen outperforms CoPilot in awareness, correctness, and coverage. $A^{3}$-CodGen helps the model to be aware of all pre-installed libraries, whereas GitHub CoPilot randomly uses third-party libraries. While CoPilot excels in local information reuse, it lacks in global functions and third-party libraries reuse.
Our method has potential to enhance Copilot by using knowledge retrieval to fetch relevant global functions and library information from the repository, enabling repository-aware code generation. For instance, Copilot’s limitations can be mitigated by incorporating interpreter information from the compiler to help Copilot become aware of third-party libraries. Additionally, using global function retrieval methods can assist Copilot in reusing functions from other code files, not just from neighboring tabs. Structuring the information before passing it to Copilot, rather than providing all information together, may also improve performance.
Finding 5: GitHub CoPilot excels in local information reuse, but its global and third-party library reuse is suboptimal. $A^{3}$-CodGen performs comparably to CoPilot in local reuse while significantly outperforming in global and third-party library reuse due to structured knowledge retrieval. Our $A^{3}$-CodGen can potentially enhance AI-assistant tools like GitHub CoPilot for better code generation.
@@ VI-A New Paradigm for Code Generation Tasks
Code generation tasks have traditionally focused on automating code production for specific requirements, with advancements like GitHub CoPilot and ChatGPT mastering this task. However, real-world code generation extends beyond meeting requirements; it requires generating code aligned with project-specific needs and leveraging existing knowledge for improved readability, maintainability, and practicability. Therefore, we hope this work sheds a light on this emerging but significant code generation topic: generating context-aware code that is compatible with the working environment.
VI-B Knowledge Engineering for Code Generation
@@
LLMs are sensitive to prompts and well-designed prompts are decisive in obtaining high-quality resultsĀ [80, 81, 82]. Either unorganized or too lengthy prompts are impractical. Instead, it is advisable to extract semantic-enriched knowledge from the code repository to guide the model to directional generate code. However, how to selectively fuse the different types of knowledge presents a challenge, potentially impacting the model’s awareness of different knowledge types. In SectionĀ V-A3, we find that providing diverse information does not necessarily improve model performance. Meanwhile, in sectionĀ V-D, we find that providing the model with global awareness and third-party library information unexpectedly enhances its ability to reuse local functions. This underscores the interaction of diverse code knowledge types and stresses the significance of efficiently mining and utilizing code knowledge in future studies.
VI-C Dataset Construction for Context Aware Code Generation
VI-C1 Benchmark Representativeness
Experiments use 29 meticulously selected Python code repositories, emphasizing high quality indicated by stars and branches, while avoiding data leaks. These repositories collectively represent the Python programming ecosystem. Labeling all functions is highly labor-intensive. Therefore, as described in SectionĀ IV-B2, we perform sampling to ensure the estimated accuracy within the sample remains within a 0.05 error range while including functions from each code repository to ensure representativeness, with an average of 13.2 functions per repository.
VI-C2 Benchmark Construction for Different Contexts
In this paper, we validate the model’s awareness of local functions, global functions, and third-party libraries. Future studies could include a more diverse range of repositories to test the framework’s effectiveness in different contexts, such as awareness of classes, inherited functions, and variables. The dataset construction method can follow the approach in SectionĀ IV-B2, where we first label functions to create a benchmark, then evaluate the generated code by predicting labels, and finally calculate reuse awareness and correctness.
In the future, we will maintain and update this dataset, providing the community with a valuable resource for evaluating repository-aware code generation.
VI-D Threats to Validity
The threats to validity of this paper lies in the following aspects.
VI-D1 Subjective Labeling
Our study involves manual labeling, introducing subjectivity in annotations. We mitigate this by engaging two professional developers for labeling, resolving any instances with ambiguous labels to establish a reliable ground truth. The dataset is public for community validationĀ [19].
VI-D2 Library Dependencies Control
Due to the complexity of library dependencies, constructing all 29 repository library dependencies is challenging. Thus, we do not consider the library version for $A^{3}$ code generation in our experiments. However, users can feasibly extract third-party libraries and generate code using the ā€˜pip list’ command or by checking the Python Interpreter in their environment to specify specific versions of libraries, as mentioned in SectionĀ III-A2.
VI-D3 Generalizing the A3-CodGen Framework to Other Programming Languages
While the experiments in this paper are conducted on Python code repositories, our $A^{3}$-CodGen framework is not limited to Python. To adapt it to another language, such as Java, developers need to use language-specific program analysis tools (e.g., SpoonĀ [83] for Java) for the Repository Knowledge Base Construction. This involves extracting relevant repository knowledge, such as signatures and source code. Other steps, including repository-aware knowledge retrieval and $A^{3}$ code generation, remain unchanged. Although constructing the knowledge base may require additional engineering effort, it does not involve new scientific innovations. It simply requires following our outlined approach using the appropriate program analysis tools for the specific language.